A grand father clock ticks each time the pendulum passes through the lowest point .if the pendulum is modeled as a simple pendulum .hiw long must it be for the ticks to occur once a second?

1
Expert's answer
2021-12-29T12:34:00-0500

since clock ticks each time the pendulum passes through the lowest point,

period of pendulum:


"T=2\pi \sqrt{L/g}=2" s


pendulum length:


"L=\frac{gT^2}{4\pi^2}=\frac{g}{\pi^2}=0.994" m